OpenStudy (anonymous):

let z=5(cos35 degrees + isin35 degrees) and w=4(cos15 degrees + isin15 degrees). a) find zw. write answer in standard form b)find z/w. write answer in trig form

OpenStudy (anonymous):

\[ z= 5 e^{ 35 i}\\ w=4 e^{ 15 i}\\ z w= 20 e^{(35+15) i}=20 e^{50 i}\\ \frac z w=\frac 5 4 e^{(35-15)i}=\frac 5 4 e^{20i} \]You can finish it now.

OpenStudy (anonymous):

Thanks so much!! This helps me understand the rest of my homework now!

OpenStudy (anonymous):

YW